OpenStudy (anonymous):

A 10.0 kg ball falling at 10.0 m/s hits a pile of feathers and comes to a complete stop in 10 s. What force does the pile of feathers exert to stop the ball? 10 N 20 N 100 N 200 N I think the answer is 100 N

OpenStudy (anonymous):

the formula for force is F=MA ( mass x acceleration)

OpenStudy (anonymous):

so i did 10.0 and 10.0 but idk what to do about the 10s

OpenStudy (whpalmer4):

The ball was going 10.0 m/s, and the pile of feathers slowed it to 0.0 m/s, taking 10 seconds to do so. You calculate the acceleration from that...

OpenStudy (whpalmer4):

Obviously, the acceleration depends on the time, right? A car that accelerates to 60 mph in 6 seconds has more powerful acceleration than one that takes 6 minutes :-)

OpenStudy (anonymous):

ah okay this makes it much simpler

OpenStudy (anonymous):

use the same formula one for acceleration?

OpenStudy (anonymous):

*or the one for acceleration

OpenStudy (whpalmer4):

well, you are accelerating the ball from 10.0 m/s to 0 m/s in 10 seconds. \(\Delta v = -10.0 \text{ m/s}\) and \(t = 10 \text{s}\)

OpenStudy (whpalmer4):

you need to find \(a\). Do you have a formula for that?

OpenStudy (anonymous):

No

OpenStudy (whpalmer4):

No? What about \(a = \Delta v/\Delta t\) (\(\Delta t = 10\text{ s}\))

OpenStudy (anonymous):

-100 is what I got

OpenStudy (whpalmer4):

the units will be (m/s)/s = m/s^2 which are the dimensions of acceleration

OpenStudy (anonymous):

ah alright

OpenStudy (whpalmer4):

-10/10 = -100?

OpenStudy (anonymous):

sorry I accidentally multiplied

OpenStudy (anonymous):

-1

OpenStudy (whpalmer4):

I get -1 m/s^2 for the acceleration. (we can ignore the - sign, that just indicates that it is slowing the ball) If m = 10.0 kg and a = -1 m/s^2, what is F?

OpenStudy (anonymous):

10

OpenStudy (whpalmer4):

Right.

OpenStudy (anonymous):

thank you I asked that question in physics and got the wrong answer so thank you

OpenStudy (whpalmer4):

Yeah, I don't think the other guy's answer is correct, as it doesn't take the time of deceleration into account

OpenStudy (anonymous):

lol why are you on mathematics your skill belongs in the science area

OpenStudy (anonymous):

you're awesome